OpenStudy (anonymous):

What is the Length of a segment whose endpoints are (3,-2) and (7,6)?

hartnn (hartnn):

Length of segment with end-points x1,y1 and x2,y2 is given as \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

OpenStudy (muhammad_nauman_umair):

u can use distance formula

OpenStudy (anonymous):

wats the actually answer

hartnn (hartnn):

did you try to first find it by yourself ?

OpenStudy (anonymous):

yes i didnt get it im confused

hartnn (hartnn):

can you show what you have tried ? we will try to spot if there is any error in your working :)

OpenStudy (anonymous):

3-2=1 7-6= 1

hartnn (hartnn):

we need to find differences in corresponding x co-ordinate so find these instead 7-3 =...? 6- (-2) =... ?

OpenStudy (muhammad_nauman_umair):

sqrt(80)

OpenStudy (anonymous):

6-(-2)=8 7-3=4

hartnn (hartnn):

yes, now plug those in the formula \(d= \sqrt{8^2 +4^2 } =... ?\)

OpenStudy (anonymous):

=80

hartnn (hartnn):

that 80 is under square root sign, right ? so your length is indeed \(\sqrt{80}\)

OpenStudy (anonymous):

the options are 2 21 2 29 4 3 4 5 sorry in between the spaces are square roots

hartnn (hartnn):

ok, so can you factor 80 ?

OpenStudy (anonymous):

the square root of 80 is is

OpenStudy (anonymous):

16

hartnn (hartnn):

\(80 = 16\times 5 \\ \sqrt {18} = \sqrt{16}\times \sqrt5\) and you know whats the square root of 16 ?

OpenStudy (anonymous):

ok got tht now wat

hartnn (hartnn):

what did u get as your final answer ?

OpenStudy (anonymous):

idk can u just plz tell me the final answer

hartnn (hartnn):

whats the square root of 16 ?

OpenStudy (anonymous):

is it 4

hartnn (hartnn):

yes! so \(\sqrt {80} =4\sqrt 5\) :)

OpenStudy (anonymous):

so then wat

hartnn (hartnn):

4 sqrt 5 is one of your option!